OpenStudy (anonymous):

m= 4/3 ; (0,5/9) write equation of line in slope intercept form

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ % (a,b) &({\color{red}{ 0}}\quad ,&{\color{blue}{ \frac{5}{9}}})\quad \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{4}{3} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

slope intercept form is y = mx + b slope(m) = 4/3 (0,5/9)....x = 0 and y = 5/9 now we sub and since we already have slope(m) and our points (x,y), we basically need to find the y intercept (b) y = mx + b 5/9 = 4/3(0) + b 5/9 = b so our equation in slope intercept form is : y = 4/3x + 5/9